OpenStudy (anonymous):

Given the function f(x; y) = 2x^2 + y + 3y^2 -1 and the unit vector u =<1/sqrt(5),2/sqrt(5)> find all points(x,y) such that the directional derivative in direction of u is 1.

OpenStudy (anonymous):

simplifying, \[ \nabla_\vec{u}f(x,y)={\partial\over\partial x}f(x,y)\times{1\over\sqrt{5}}+{\partial\over\partial y}f(x,y)\times{2\over\sqrt{5}} \]

OpenStudy (anonymous):

\[ \nabla_\vec{u}f(x,y)={1\over\sqrt{5}}\left[(4x)+2(1+6y)\right] \]

OpenStudy (anonymous):

yes that is right. the problem is if i expand this equation and eqaute it to 1, i get a function of two variables x and y which is (4x+12y+2)^2=5. I cant get the other equation for me to solve simultaneously.

OpenStudy (anonymous):

you'd get a "locus" of points...

OpenStudy (anonymous):

probably a circle or an ellipse

OpenStudy (anonymous):

ok, I guess you are right there thanks.